Using Rodrigue’s formula, obtain expression for the Hermite polynomial
H3(x)
and show that
H3'(x) = 6H2(x)
1
Expert's answer
2020-02-26T10:16:01-0500

The Rodrigues formula for the Hermite Polynomial can be written as

"H_n(x)=(-1)^xe^{x^2}\frac{d^n}{dx^n}e^{-x^2}"

For "H_2(x)" we obtain

"H_2(x)=(-1)^xe^{x^2}\frac{d^2}{dx^2}e^{-x^2}=4x^2-2"

For "H_3(x)"

"H_3(x)=(-1)^xe^{x^2}\frac{d^3}{dx^3}e^{-x^2}=8x^3-12x"

"H'_3(x)=24x^2-12=6(4x^2-2)=6H_2(x)"